Your boss Sally Maloney, treasuer of Fred Clark Enterprises aske you to help her estimate the intrinsic value of the company's stock.
1
Expert's answer
2012-07-26T11:25:11-0400
Intrinsic value indicates the true value of a particular business based on ephemeral factors like research, brand identity and market psychology.
Before beginning the calculation, you will need to have a specific investment time horizon in mind and be able to estimate the future earnings per share over a specific time horizon.
Here are some assumptions for company x.
You have an investment horizon of 10 years.
The company's EPS (Earnings Per Share) is currently $2.50.
You estimate that the company's EPS will grow at a steady rate of 10% per year over the next 10 years.
The company's average PE ratio has been around 15 over the past several years, and you expect this trend to continue into the forseeable future.
The company has an average dividend payout of 3%.You would like the stock to return at least 11 percent per year.
Using this information, there is a basic formula for calculation the intrinsic value of company x's stock:
Forecasted Stock Price in 2020 = Earnings Per Share after the 10th year X Average PE
Ratio

Earnings Per Share After the 10th Year = Current EPS X Rate of EPS Increase , in this case…
Earnings Per Share After the 10th Year = 2.50 X (1.10 ^10) = $6.48
Forecasted Stock Price in 2020 = $6.48 X 15 = $97.20

Next, we also need to calculate how much the dividends will be worth. Dividend Payout = Total Dividends / Total Earnings Per Share WHERE Total Dividends = Total EPS X Average
Dividend Payout

Total Dividends = $43.83 X 3.0% = $1.32

Net Present Value = Future Value / Expected ROI = $98.52 / (1.10 ^ 10) = $98.52 / 2.59 = $37.98

Since the calculated intrinsic value of $37.98 is more than the current share price of $30.00, you should just go ahead and buy the
stock.